# ReelWave: A Multi-Agent Framework Toward Professional Movie Sound Generation ![[Uncaptioned image]](x1.png)

### v1

@@
Abstract
Film production is an important application of generative audio, where richer context is provided through multiple scenes. In ReelWave, we propose a multi-agent framework for audio generation inspired by the professional movie production process. We first capture semantically and temporally synchronized ‘on-screen’ sound by training a prediction model that forecasts three interpretable, time-varying audio control signals: loudness, pitch, and timbre. These three parameters are then specified as conditions by a cross-attention module. Our framework subsequently infers ‘off-screen’ sound to complement the generation through cooperative interaction between communicative agents. Each agent takes on specific roles similar to those of a movie production team and is supervised by an agent called the Sound Director. Furthermore, we explore the case when the conditional video consists of multiple scenes, a scenario commonly seen in videos extracted from movies of considerable length. As a result, our framework can capture a richer context for audio generation conditioned on video clips extracted from movies.
1 Introduction
@@
While generating audio from text has been an emerging research topic in recent years, there has been a shift toward video-conditioned audio generation. Much of the work in this area has focused on aligning video and audio modalities. However, even state-of-the-art synchronized audio still falls short of meeting the demands of real-world sound production in the TV and movie industry, see Figure ReelWave: A Multi-Agent Framework Toward Professional Movie Sound Generation. For more immersive visual storytelling, sound design goes beyond synchronized sound effects to include music and dialogue. ‘On-screen’ sound effects, or Foley sounds, are those that need precise synchronization with the visual action. In contrast, ‘off-screen’ sound contributes to the overall context of the scene, such as dialogue, environment sounds or background music. Additionally, generating visual sound without recognizing multiple scenes often leads to suboptimal performance, as noted in [38], failing to meet the increasing demand for generative audio in longer, multi-scene videos.
@@
Recent advances in large language models have made it possible to use LLM for task planning and carrying out a variety of downstream tasks, including the planning and execution of robot tasks and general human motion generation [31]. Confronted with the complexities of real-world tasks, works such as CAMEL [16] proposed a communicative agent framework that performs task planning with multiple rounds of conversation. Although audio generation can potentially involve complexity in planning and generation, particularly when the condition incorporates multiple events in text or multiple scenes in video, most of the current work only focuses on single-round inference, with few involving LLM in explicit planning and none in providing feedback and revision in audio generation.
@@
In this paper, we introduce ReelWave, a multi-agent framework for movie sound generation. Our framework draws inspiration from the real filmmaking process. We first use a training-free scene detector to divide the conditional video into multiple scenes, then utilize a video understanding model to obtain the textual description of the visual scene. To produce aligned on-screen sound, we train a prediction model to predict three interpretable time-varying audio control signals of the to-be-generated sound from video, namely loudness, pitch, and timbre. To complement the generation with off-screen sound, we utilize a role-playing setting that incorporates a sound director who supervises the generation process, a voice actor for dialogues, a composer to compose music, and finally, a mixer to mix multiple generated soundtracks with the correct sequence and reasonable volume. Our contributions are summarized as:
@@
To the best of our knowledge, our framework is among the first to incorporate a multi-agent framework for multi-scene audio generation. We believe ReelWave will contribute to sound design artists and editors by providing a more expressive and interpretable generation tool. We highly encourage the reader to listen to our audio examples, which demonstrate the efficacy of ReelWave.
2 Related Work
@@
LLM-based Agent Method. Recent advancements in large language models (LLMs) have fostered significant progress in using LLMs for task planning and execution across various domains, such as robot task planning and software development [5, 25, 36]. LLMs demonstrate the ability for zero-shot or few-shot generalization, enabling efficient transfer of tasks without the need for significant changes in the model parameters [32]. In the domain of audio generation, WavJourney [22] has pioneered the use of LLMs to connect different audio models for creative audio production.
@@
To address this, CAMEL [16] explores the use of a multi-agent framework to divide and conquer tasks, with each agent adopting the role of either an AI Assistant or AI User. This method improves scalability and ensures a comprehensive approach to problem-solving, as each agent addresses the task from different perspectives. In our work, we extend this framework by introducing a dedicated Sound Director agent. The Sound Director supervises the generation of the audio components (music, dialogue, soundtrack) in a collaborative manner, providing feedback across multiple rounds if the initial plans prove unsatisfactory. This mirrors the real-world process of movie sound production, where the sound director continuously works with the sound team to adjust the generated audio based on evolving creative needs.
@@
Diffusion-based Audio Generation. The introduction of the latent diffusion method for audio generation was a significant breakthrough in the field, with works like AudioLDM [19] leading the way. Subsequent models, such as Tango [8], AudioLDM-2 [20], and Auffusion [35], have used pre-trained LLMs such as Flan-T5 [3] for efficient text encoding, demonstrating high-quality results through large-scale training. These text-conditioned methods have opened the door to using visual input for audio generation, offering a more efficient alternative to training video-to-audio generators from scratch by augmenting existing audio generation frameworks with visual conditioning [29].
@@
FoleyCrafter [40] pushes the boundaries by incorporating additional modules to adapt visual semantics, detect timestamps, and align temporal conditions for generating audio. ReWaS [14] further advances this by integrating continuous timestamp conditions, though it still lacks consideration of visual semantic style and relies mainly on input captions. Our framework builds upon these approaches by introducing a comprehensive multi-agent setup to model both the ”on-screen” and ”off-screen” audio conditions, enabling a more nuanced and coherent audio generation process for longer video segments. In particular, our approach allows for the exploration of multiple scenes, which is a common characteristic of real-world movie productions.
@@
Integration of Visual and Temporal Conditions for Audio Generation. Text-to-audio (TTA) generation has seen significant progress with models like Make-an-Audio 2 [10], which was among the first to handle complex textual descriptions for audio generation. However, it primarily focuses on the sequencing of generated audio and lacks the flexibility for multi-turn editing. Additionally, Make-an-Audio 2 is limited by extensive training with LLM-based data augmentation, restricting its applicability to TTA tasks. AudioGPT [11], WavJourney [21], and WavCraft [18] represent further advances in compositional audio generation. However, these systems still rely on off-the-shelf generators and do not leverage the full potential of GPT-4’s capabilities. Their text-based descriptions often result in audio that doesn’t meet user-specified length or complexity requirements.
@@
For video-to-audio (VTA) generation, works such as FoleyGen [24] and V2meow [27] have employed transformers to model the audio generation process. However, these autoregressive models tend to suffer from accumulating errors during inference, which can degrade the quality of long-form audio generation [17]. Our method addresses this limitation by using a multi-agent framework, which provides iterative feedback and enhances the stability of long-form, multi-scene audio generation, ensuring higher consistency and quality over time.
@@
Comparison to Previous Work Our approach offers several advantages over previous methods. Firstly, the incorporation of a multi-agent framework with a dedicated Sound Director allows for iterative refinement and collaboration throughout the audio generation process. This is a key feature that distinguishes our work from methods that rely on a single-stage generation process or those lacking feedback mechanisms. Secondly, our framework is designed for long-form, multi-scene audio generation, enabling it to handle more complex and context-rich audio tasks that are typical in movie production. Lastly, by leveraging both visual and temporal conditions in a unified model, we ensure that our generated audio is not only contextually relevant but also richly detailed, matching the demands of professional movie sound design.
3 Method
@@
The on-screen sound is produced by the Foley Artist, which encompasses a sound prediction module given the video condition. To further enrich the generation content, additional off-screen sound is produced by the three agents presented in the middle of Figure 2. Considering single agent planning framework may have limited perspective and thus a higher error rate, we utilize a supervisor agent to offer critique and guidance during the planning process. In addition to generating the accompanying audio from video clips, ReelWave may take a text-only plot summary from the user and generate multi-scene sound, denoted as the pink dash line in Figure 2. Through generating multiple overlapping soundtracks, we take a step further to the real-world movie sound generation.
3.1 Scene detection and understanding
V2A-SceneDetector proposed an efficient way to recognize scene boundary within a given video by computing the similarity difference between consecutive frames. However, the authors did not fully demonstrate the potential of this method by limiting the experiments on the VGGSound dataset, where the majority are single-scene videos. Our framework employs this scene detector to partition a given movie-like video input into segments. Afterward, we use a video understanding model to capture the key event description and full description.
The key event description reflects simple and atomic auditory elements encompassed in the video, such as the type of sound or the event causing it. For example, “The sound of explosions” and “The sound of roaring engines”. We use this concise description as the caption for on-screen sound production. The full description aims to capture the full ongoing event and visual environment, which provides the context for the framework to infer off-screen sound.
3.2 Multi-agent generation planning
Script Writer. The full description of the video is passed to the Script Writer to convert into a structured script, similar to the storyboard used in the movie industry. The Script Writer aims to infer the storyline behind the visual scene, in addition to the environment setting suitable for the story to happen. The script consists of five key elements:
@@ 3.2 Multi-agent generation planning
Foley Artist. Foley Artist consists of two parts: on-screen Foley sound generator and off-screen Foley sound planner. We define on-screen sound as the main events that happen on the screen. The on-screen Foley sound is generated using the key event description from a video understanding model, with Loudness, Pitch, Timbre, as the conditions for producing synchronized sound. Further details are illustrated in Section 3.3.
@@
Off-screen Foley sound corresponds to environment sound that illuminates the event’s ambiance, such as “distant screams of townspeople” or “occasional distant gunfire” in a war scene. The script is provided to the off-screen Foley sound planner. To better complement the on-screen Foley sound, we provide the key event description together with the script. The agent is asked to focus on generating ambient or event sounds while excluding human voices and music. In addition, the description for each soundtrack should be simple and atomic, directly describing the type of sound or the event making the sound. The output of the sound planner is a detailed Foley design plan for off-screen sound generation. See supplementary material for sample JSON Foley plan generated.
@@
Composer. The task for the Composer is to translate the audio script into a detailed music design plan for generation (see supplementary material for music plan). The agent is asked to focus on generating only the music, with the description capturing the style and instruments for each soundtrack. If the agent thinks music is unnecessary, it will respond N/A. Only the script is provided for reference.
@@
Voice Actor. The task for Voice Actor is to translate the audio script into a detailed voice design plan for generation. (see supplementary material for voice plan). The agent focuses only on generating the human voice, with one dialogue for each soundtrack. To stylize the generation, either a male voice or a female voice can be specified. We additionally add a total word constraint for every dialogue, assuming two words for one second, which is close to human speaking speed in normal conversation. Similar to the Composer agent, only the script is provided, with the agent responding N/A if the scene does not have human dialogues.
@@
Mixer. After obtaining the textual plan from the Foley Artist, Composer, and Voice Actor, the Mixer will conduct a final mix of multiple soundtracks to ensure the right balancing and time order (see supplementary material for mixer plan). Specifically, it will assign appropriate volume and adjust the start time and end time to each soundtrack. The volume is an integer value in dB following LUFS standard. For background sound effect/music, the usual level is around to dB. For speech is around dB.
@@
Sound Director. The sound director conducts a two-way conversation with the Foley Artist, Composer and Voice Actor, providing feedback and instructions to refine the generation plan. In case any inconsistency happens between the agent’s action and the task, the director will provide feedback to rectify the agent’s behavior. This is similar to the movie industry, where the soundtracks are not finalized in one go but rather through multiple rounds of refinement between the production team and the director. In our framework, we conduct multiple rounds of the conversation between individual generation agents and the director. The conversation ceased when the director output token. We prompt the director to enrich the design plan with more details and instruct the planning agent to add additional soundtracks when appropriate. In the case when the input is plot summary, the director may ask the Script Writer to add additional scenes if necessary.
@@
Output structure. All communication is operated through text, and the agents will output a plan in JSON format; see supplementary material. The JSON file is further used for the generation of backbones to generate audio soundtracks. Our design offers a transparent, structured way for movie sound generation that is easy to monitor. In addition, the user may optionally provide the custom dialogue lines to suit his or her need. Please see the supplementary for JSON examples for Foley plan, music plan and voice plan.
3.3 Audio-conditioning for on-screen sound

### v2

# ReelWave: Multi-Agentic Movie Sound Generation through Multimodal LLM Conversation

@@
Abstract.
Current audio generation conditioned by text or video focuses on aligning audio with text/video modalities. Despite excellent alignment results, these multimodal frameworks still cannot be directly applied to compelling movie storytelling involving multiple scenes, where ‘on-screen’ sounds require temporally-aligned audio generation, while ‘off-screen’ sounds contribute to appropriate environment sounds accompanied by background music when applicable. Inspired by professional movie production, this paper proposes a multi-agentic framework for audio generation supervised by an autonomous Sound Director agent, engaging multi-turn conversations with other agents for on-screen and off-screen sound generation through multimodal LLM. To address on-screen sound generation, after detecting any talking humans in videos, we capture semantically and temporally synchronized sound by training a prediction model that forecasts interpretable, time-varying audio control signals: loudness, pitch, and timbre, which are used by a Foley Artist agent to condition a cross-attention module in the sound generation. The Foley Artist works cooperatively with the Composer and Voice Actor agents, and together they autonomously generate off-screen sound to complement the overall production. Each agent takes on specific roles similar to those of a movie production team. To temporally ground audio language models, in ReelWave, text/video conditions are decomposed into atomic, specific sound generation instructions synchronized with visuals when applicable. Consequently, our framework can generate rich and relevant audio content conditioned on video clips extracted from movies.
1. Introduction
@@
Following text-guided audio generation, more research attention has been paid to video-conditioned audio generation, specifically, on aligning video and audio modalities. Despite excellent results, state-of-the-art synchronized audio still falls short of meeting the requirement for real-world sound production for video clips much less for movies and films, for instance. For more immersive visual storytelling, sound design goes beyond synchronized sound effects to include music and dialogue. ‘On-screen’ sound effects, either human speech or Foley sounds, are those that need precise synchronization with the visual action. On the other hand, ‘off-screen’ sound contributes to the overall context of the scene, such as dialogue, environment sounds, or background music. Additionally, generating visual sound without recognizing multiple scenes often leads to suboptimal performance, as noted in (Yi and Li, 2024), failing to meet the increasing demand for generative audio in longer, multi-scene videos.
@@
Recent advances in large language models have made it possible to use LLM for task planning and carrying out a variety of downstream tasks, including the planning and execution of robot tasks and general human motion generation. Confronted with the complexities of real-world tasks, works such as CAMEL (Li et al., 2023) proposed a communicative agent framework that performs task planning with multiple rounds of conversation. Although audio generation can potentially involve complexity in planning and generation, particularly when the condition incorporates multiple events in text or multiple scenes in video, most of the current work only focuses on single-round inference, with few involving LLM in explicit planning and none in providing feedback and revision in audio generation.
@@
In this paper, we introduce ReelWave, a multi-agent framework for movie sound generation. As shown in Figure 1, our framework draws inspiration from the real filmmaking process. We first use a training-free scene detector to divide the conditional video into multiple scenes, then utilize a video understanding model to obtain the textual description of the visual scene. To produce aligned on-screen sound, we either utilize a lip-to-speech model to generate speech or a prediction model to predict three interpretable time-varying audio control signals of the to-be-generated sound from video, namely loudness, pitch, and timbre. To complement the generation with off-screen sound, we utilize a role-playing setting that incorporates a sound director who supervises the generation process, a voice actor for dialogues, a composer to compose music, and finally, a mixer to mix multiple generated soundtracks with the correct sequence and reasonable volume. Our contributions are summarized as:
@@
To the best of our knowledge, our framework is among the first to introduce multiple, autonomous sound making agents operating in a multimodal conversational LLM framework for multi-scene audio generation. We believe ReelWave will contribute to sound design artists and editors by providing a more expressive and interpretable generation tool. We highly encourage the reader to view and listen to our audio examples in the supplemental video, Figure 2.
2. Related Work
@@
LLM-based Agent Method. Recent advancements in large language models (LLMs) have fostered significant progress in using LLMs for task planning and execution across various domains, such as robot task planning and software development (Driess et al., 2023; Rawles et al., 2024; Yang et al., 2023). LLMs demonstrate the ability for zero-shot or few-shot generalization, enabling efficient transfer of tasks without the need for significant changes in the model parameters (Xi et al., 2023). In the domain of audio generation, WavJourney (Liu et al., 2023b) and WavCraft (Liang et al., 2024) pioneered the use of LLMs to connect different audio models for creative audio production. However, these systems are limited to a single-agent setting, which may fall short when the task becomes more complex and do not leverage the full potential of GPT-4’s capabilities, often resulting in audio that doesn’t meet user-specified length or complexity requirements.
@@
To enhance system capacity, CAMEL (Li et al., 2023) explores the use of a multi-agent framework to divide and conquer tasks related to math, coding, and science, with each agent adopting the role of either an AI Assistant or AI User. This method improves scalability and ensures the conversation follows the right direction. Inspired by this work, we introduce a supervisor agent over multiple planning and generation agents. The supervisor agent coordinates the generation of the audio components (music, dialogue, soundtrack) in a collaborative manner, providing feedback across multiple rounds if the initial plans prove unsatisfactory. This mirrors the real-world process of movie sound production, where the sound director continuously works with the sound team to adjust the generated audio based on evolving creative needs.
@@
Diffusion-based Audio Generation. The introduction of the latent diffusion method for audio generation was a significant breakthrough in the field, with works like AudioLDM (Liu et al., 2023a) leading the way. Subsequent models, such as Tango (Ghosal et al., 2023), AudioLDM-2 (Liu et al., 2024), and Auffusion (Xue et al., 2024), have used pre-trained LLMs such as Flan-T5 (Chung et al., 2022) for efficient text encoding, demonstrating high-quality results through large-scale training. These text-conditioned methods have opened the door to using visual input for audio generation, offering a more efficient alternative to training video-to-audio generators from scratch by augmenting existing audio generation frameworks with visual conditioning (Wang et al., 2024).
@@
FoleyCrafter (Zhang et al., 2024a) pushes the boundaries by incorporating additional modules to adapt visual semantics, detect timestamps, and align temporal conditions for generating audio. ReWaS (Jeong et al., 2024) further advances this by integrating continuous timestamp conditions, though it still lacks consideration of visual semantic style and relies mainly on input captions. Our multimodal LLM conversational framework builds upon these technical contributions, by introducing a comprehensive multi-agent setup to model both the “on-screen” and “off-screen” audio conditions, enabling a more nuanced and coherent audio generation process for longer video segments. In particular, our approach allows for the exploration of multiple scenes, which is a common characteristic of real-world movie productions.
@@
Integration of Visual and Temporal Conditions for Audio Generation. Text-to-audio (TTA) generation has seen significant progress with models such as Make-an-Audio 2 (Huang et al., 2023), which was among the first to handle complex textual descriptions for audio generation. However, this work primarily focuses on sequencing of generated audio and lacks the flexibility for multi-turn editing. Additionally, Make-an-Audio 2 is limited by extensive training with LLM-based data augmentation which restricts its applicability to TTA tasks.
@@
For video-to-audio (V2A) generation, works such as FoleyGen (Mei et al., 2024) and V2meow (Su et al., 2024) have employed transformers to model the audio generation process. However, these autoregressive models tend to suffer from accumulating errors during inference, which can degrade the quality of long-form audio generation (Li et al., 2024). Our method addresses this limitation by using a multi-agent framework, which provides iterative feedback and enhances the stability of long-form, multi-scene audio generation, ensuring higher consistency and quality over time.
@@
Comparison to Previous Work. Our approach offers several advantages over previous methods. Firstly, the incorporation of a multi-agent framework with a dedicated Sound Director allows for iterative refinement and collaboration throughout the audio generation process. This is a key feature that distinguishes our work from methods that rely on a single-stage generation process or those lacking feedback mechanisms. Secondly, our framework is designed for long-form, multi-scene audio generation, enabling it to handle more complex and context-rich audio tasks that are typical in movie production. Lastly, by leveraging both visual and temporal conditions in a unified model, we ensure that our generated audio is not only contextually relevant but also richly detailed, matching the demands of professional movie sound design.
The framework closest to ours is ComposerX (Deng et al., 2024), which proposes using a multi-agent framework to generate symbolic music notation for melody, harmony, and instrumentation. On the other hand, pure music generation is a far more structured and limited domain compared to general video/text-conditioned audio generation, where the latter demands temporal precision and semantic diversity. The complexities involved in aligning audio to the real-world video, understanding cross-modal relationships, and modeling a multi-agent system to generate audio pose extra challenges beyond those for generating music components.
3. Method
@@
The on-screen sound is produced either by the Foley Artist, which encompasses a sound prediction module given the video condition, or the Voice Actor, which consists of a pretrained lip-to-speech module. To further enrich the generation content, additional off-screen sound is produced by the three agents presented in the middle of Figure 1. Considering single agent planning framework may have a limited perspective and thus a higher error rate, we utilize a supervisor agent to offer critique and guidance during the planning process. In addition to generating the accompanying audio from video clips, ReelWave may take a text-only plot summary from the user and generate multi-scene sound, denoted as the pink dash line in Figure 1. Through generating multiple overlapping soundtracks, we take a step further toward real-world movie sound generation.
3.1. Scene detection and understanding
V2A-SceneDetector (Yi and Li, 2024) proposed an efficient way to recognize scene boundaries within a given video by computing the similarity difference between consecutive frames. However, the authors did not fully demonstrate the potential of this method by limiting the experiments to the VGGSound dataset, where the majority are single-scene videos. Our framework employs this scene detector to partition a given movie-like video input into segments. Let represent the sequence of frame embeddings, where is the number of frames, and is the embedding dimension. Then we compute pairwise similarities between frame embeddings followed by centering, where self-similarity matrix :
followed by computing absolute differences between consecutive rows:
and identifying significant peaks in using a pre-defined threshold:
Finally, the scene boundary is defined as
where are detected peak indices.
Afterward, we use a video understanding model (Zhang et al., 2024c) to capture the key event description and full description. The key event description reflects simple and atomic auditory elements encompassed in the video, such as the type of sound or the event causing it. For example, “The sound of explosions” and “The sound of roaring engines.” We use this concise description as the caption for on-screen sound production. The full description aims to capture the full ongoing event and visual environment, which provides the context for the framework to infer off-screen sound.
In addition, the video understanding model checks whether there exists any talking human in the input video. If so, a pretrained lip-to-speech model is used for audible voice generation. We treat this as the on-screen sound. Otherwise, we proceed by passing the key event description to the Foley Artist for on-screen sound generation. Details are described in Section 3.2.
3.2. Multi-agentic generation and planning
Script Writer. The full description of the video is passed to the Script Writer to convert into a structured script, similar to the storyboard used in the movie industry. The Script Writer aims to infer the storyline behind the visual scene, in addition to the environmental setting suitable for the story to happen. The script consists of five key elements:
@@ 3.2. Multi-agentic generation and planning
Foley Artist. Foley Artist consists of two parts: on-screen Foley sound generator and off-screen Foley sound planner. We define on-screen Foley sound as the main events that happen on the screen, which is generated using the key event description from the video understanding model, with Loudness, Pitch, Timbre, as the conditions for producing synchronized sound. Further details are illustrated in Section 3.3.
@@
Off-screen Foley sound corresponds to environmental sound that illuminates the event’s ambiance, such as “distant screams of townspeople” or “occasional distant gunfire” in a war scene. The script is provided to the off-screen Foley sound planner. To better complement the on-screen Foley sound, we provide the key event description together with the script. The agent is asked to focus on generating ambient or event sounds while excluding human voices and music. In addition, the description for each soundtrack should be simple and atomic, directly describing the type of sound or the event making the sound. The output of the sound planner is a detailed Foley design plan for off-screen sound generation. See Figure 11 for sample JSON Foley plan generated.
@@
Composer. The task for the Composer is to translate the audio script into a detailed music design plan for generation (see Figure 12 in supplementary material for music plan). The agent is asked to focus on generating only the music, with the description capturing the style and instruments for each soundtrack. If the agent thinks music is unnecessary, it will respond N/A. Only the script is provided for reference.
@@
Voice Actor. For off-screen voice generation, the Voice Actor will translate the audio script into a detailed voice design plan for generation. See Figure 9 in supplementary material for voice plan. The agent focuses only on generating the human voice, with one dialogue for each soundtrack. To stylize the generation, either a male voice or a female voice can be specified. We additionally add a total word constraint for every dialogue, assuming two words per second, which is close to human speaking speed in normal conversation. Similar to the Composer agent, only the script is provided, with the agent responding N/A if the scene does not have human dialogues.
For on-screen voice generation, in the case when the video understanding model detects a person talking, the agent will use a pretrained lip-to-speech model to directly generate human voice from input silent video with lip synchronization.
@@
Mixer. After obtaining the textual plan from the Foley Artist, Composer, and Voice Actor, the Mixer will conduct a final mix of multiple soundtracks to ensure the right balancing and time order (see Figure 10 for mixer plan). Specifically, it will assign appropriate volume and adjust the start time and end time to each soundtrack. The volume is an integer value in dB following the LUFS standard. For background sound effect/music, the usual level is around to dB. For speech is around dB.
@@
Sound Director. The sound director conducts a two-way conversation with the Foley Artist, Composer and Voice Actor, providing feedback and instructions to refine the generation plan, Figure 4. In case any inconsistency happens between the agent’s action and the task, the director will provide feedback to rectify the agent’s behavior. This is similar to the movie industry, where the soundtracks are not finalized in one go but rather through multiple rounds of refinement between the production team and the director. In our framework, we conduct multiple rounds of conversation between individual generation agents and the director. The conversation ceased when the director output token. We prompt the director to enrich the design plan with more details and instruct the planning agent to add additional soundtracks when appropriate. In the case when the input is a plot summary, the director may ask the Script Writer to add additional scenes if necessary.
@@
Output structure. All communication is operated through text, and the agents will output a plan in JSON format; see supplementary material. The JSON file is further used for the generation of backbones to generate audio soundtracks. Our design offers a transparent, structured way for movie sound generation that is easy to monitor. In addition, the user may optionally provide the custom dialogue lines to suit his or her needs. Please see the supplementary for JSON examples for Foley plan, music plan, and voice plan.
3.3. Audio control signals prediction from video